# RestoreVAR: Visual Autoregressive Generation for All-in-One Image Restoration

### v1

@@ 5 Conclusions
We proposed RestoreVAR, a fast and effective generative approach for All-in-One image Restoration (AiOR). Built on the Visual Autoregressive Modeling (VAR) backbone, RestoreVAR benefits from VAR’s strong generative priors and significantly faster inference compared to latent diffusion models (LDMs). To tailor VAR for AiOR, we introduced cross-attention mechanisms that inject semantic information from the degraded image into the generation process. Additionally, we proposed a non-generative latent refiner transformer to convert discrete latents to continuous ones, along with a VAE decoder fine-tuned on continuous latents, which together improve reconstruction fidelity. RestoreVAR achieves state-of-the-art performance among generative AiOR models, significantly outperforming LDM-based methods while delivering over 10×10$\times$10 × faster inference and strong generalization to real-world degradations.

### v2

@@ 5 Conclusions
We proposed RestoreVAR, a fast and effective generative approach for AiOR. Built on the VAR backbone, RestoreVAR benefits from VAR’s strong generative priors and significantly faster inference compared to LDMs. To tailor VAR for AiOR, we introduced cross-attention mechanisms that inject semantic information from the degraded image into the generation process. Additionally, we proposed a non-generative latent refiner transformer to convert discrete latents to continuous ones, along with a VAE decoder fine-tuned on continuous latents, which together improve reconstruction fidelity. RestoreVAR achieves state-of-the-art performance among generative AiOR models, outperforming LDM-based methods while delivering over $10\times$ faster inference and strong generalization.